40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 62 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  13765   Thu Apr 19 00:03:51 2018 gautamUpdateIOOMore IMC NBing

Summary:

As shown in the Attachments, it seems like IMC DAC and coil driver noise is the dominant noise source above 30Hz. If we assume the region around the bounce peak is real motion of the stack (to be confirmed with accelerometer data soon), this NB is starting to add up. Much checking to be done, and I'd also like to get a cleaner measurement of coil driver and DAC noise for all 3 optics, as there seems to be a factor of ~5 disagreement between the MC3 coil driver noise measurement and my previous foray into this subject. The measurement needs to be refined a little, but I think the conclusion holds.

Details:

  1. I had a measurement of the MC3 coil driver noise from ~2weeks ago when I was last working on this that I had not yet added to the NB.
  2. Today I added it. To convert from measured voltage noise to frequency noise, I assumed the usual 0.016N/A per coil number, which is probably a large source of systematic error.
  3. I define the "nominal" IMC operating condition as MC1 and MC3 having the analog de-whitening filters switched on, but MC2 switched off.
  4. So length noise should be dominated by coil driver noise on MC1 and MC3, and DAC noise on MC2.
  5. The measurement I had was made with the input to the coil driver board terminated in 50ohms. Measurement was made in-situ. The measurement has a whole bunch of 60Hz harmonics (despite the Prologix box being powered by a linear power adapter, but perhaps there are other ground loops which are coupling into the measurement). So I'd like to get a cleaner measurement tmrw.
  6. To confirm, Koji suggested some On/Off test by driving some broadband noise in the coils. I figured toggling the analog de-whitening, such that the DAC noise or coil driver electronics dominate is an equally good test.
  7. Attachment #2 shows the effect in arm error and control signal spectra. Note that I engaged analog de-whitening on all 3 optics for the red curves in this plot. But even leaving MC2 de-whitening off, I could see the read curve was below the black reference trace, which was taken with de-whitening off on all 3 optics.

Remarks:

Since I sunk some time into it already, the motivation behind this work is just to try and make the IMC noise budget add up. It is not directly related to lowering the IR ALS noise, but if it is true that we are dominated by coil driver noise, we may want to consider modifying the MC coil driver electronics along with the ITM and ETMs.

  13764   Wed Apr 18 22:46:23 2018 johannesConfigurationGeneralAS port laser injection

Using Gautam's Finesse file and the cad files for the 40m optical setup I propagated the arm mode out of the AS port. For the location of the 3.04 mm waist I used the average distance to the ITMs, which is 11.321 m from the beam spot on the 2 inch mirror on the AS table close to the viewport. The 2inch lens focuses the IFO mode to a 82.6 μm waist at a distance of 81 cm, which is what we have to match the aux laser fiber output to.

I profiled the fiber output and obtained a waist of 289.4 μm at a distance of 93.3 cm from the front edge of the base of the fiber mount. Next step is to figure out the lens placement and how to merge the beam paths. We could use a simple mirror if we don't need AS110 and AS55, we could use a polarizing BS and work with s polarization, or we find a Faraday Isolator.


While doing a beam scan with the razor blade method I noticed that the aux laser has significant intensity noise. This is seen on the New Focus 1611 that is used for the beat signal between PSL and aux laser, as well as on the fiber output PD. There is a strong oscillation around 210 kHz. The oscillation frequency decreases when the output power is turned down, the noise eater has no effect. Koji suggested it could be light scattering back into the laser because I couldn't find a usable Faraday Isolator back when I installed the aux laser in the PSL enclosure. I'll have to investigate this a little further, look at the spectrum, etc. This intensity noise will appear as amplitude noise of the beat note, which worries me a little.

power_out_fluctuation_DC.png      power_out_fluctuation_AC_zoom.png

  13763   Wed Apr 18 20:33:19 2018 KevinUpdateGeneralseismometer interfaces

Steve, the pictures you posted are not the AA board I was referring to. The attached pictures show the board which is sitting beneath the GPS time server.

  13762   Wed Apr 18 19:02:15 2018 gautamUpdateIOOMC spot centering scripts

I'm working on fixing these (and the associated MEDM scripts) up so that we can get some reliable readback on how well centered the spots are on the MC mirrors. Seems like a bunch of MEDM display paths were broken, and it looks like the optimal demod phases (to put most of the output in I quadrature) are not what the current iteration of the scripts set them to be. It may well be that the gains that convert demodulated counts to mm of spot offset are also not correct anymore. I ran the script ~4times in ~1 hour time span, and got wildly different answers for the spot centering each time, so I wouldn't trust any of those numbers atm...

As you can see in Attachment #1, I stepped the demod phase of one of the servos from -180 to 180 degrees in 5degree increments. The previously used value of 57degrees is actually close to the worst possible point (if you want the signal in the I quadrature, which is what the scripts assume).

I used Attachment #2 to change up the demod phases to maximize the I signal. I chose the demod phase such that it preserved the sign of the demodulated signal (relative to the old demod phases). I also made some StripTool templates for these, and they are in the MC directory. Doing the spot centering measurement with the updated demod phases yields the following output from the script:

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[0.72506586251144134, 7.1956908534034403, 0.54754354165324848, -0.94113388241389784, -3.5325883025464959, -2.4934027726657142]

Seems totally unbelievable still that we are so far off center on MC1 yaw. Perhaps the gains and calibration to convert from counts to mm of spot offset need to be rechecked.

  13761   Wed Apr 18 17:15:35 2018 ranaConfigurationComputersNODUS: no xmgrace for dataviewer

Turns out, there is no RPM for XmGrace on Scientific Linux 7. Since this is the graphic output of dataviewer, we can't use dataviewer through X windows until this gets fixed. CDS is looking into a xmGrace replacement, but it would be better if we can hijack a alt RH repo to steal a temporary xmgrace RPM. KT has been pinged.

  13760   Wed Apr 18 16:59:35 2018 ranaUpdatePEMfinal setup sketch: EX Seis

Can you please add dimensions to the drawing, so we can see if things fit and what the cable lenghts need to be?

For the panel on the granite slab, we should use a thinner piece of metal and mount it with an offset so that the D-sub cable can be fished through the hole in the slab. The hole is wide enough for 2 cables, but not 2 connectors.


Attached is a 8-day minute trend of the heater control signals, as well as the in-loop temperature sensor (which underestimates the true fluctuations; we really need an out-of-loop sensor attached to the can or seismometer).

You can see that since the last tuning (on the 13th), its been stable at the set point of 39 C with 8.5 - 10 W of heating power. Need to add the PID loop settings (all the sliders on the MEDM screen) to the frames so that we can help in diagnosing. Also, fix the spelling of "Celcisususs".

  13759   Wed Apr 18 12:18:39 2018 KiraUpdatePEMfinal setup sketch

I've updated the sketches and added in front panels for the seismometer block and the 1U panel (attachments 3 and 4). There was an issue when it came to the panel on the block because the hole is only big enough for the cable that already exists there and there is no space to add in the D-9 connector. Not quite sure how to resolve this issue. Attachment 7 is the current panel on the seismometer block. Attachments 5 and 6 are the updated temperature circuit and the heater circuit.

The boxes will be located in the short racks at EX and EY to minimize cable length.

  13758   Wed Apr 18 10:44:45 2018 gautamUpdateCDSslow machine bootfest

All slow machines (except c1auxex) were dead today, so I had to key them all. While I was at it, I also decided to update MC autolocker screen. Kira pointed out that I needed to change the EPCIS input type (in the RTCDS model) to a "binary input", as opposed to an "analog input", which I did. Model recompilation and restart went smooth. I had to go into the epics record manually to change the two choices to "ENABLE" and "DISABLE" as opposed to the default "ON" and "OFF". Anyways, long story short, MC autolocker controls are a bit more intuitive now I think.

  13757   Tue Apr 17 14:08:29 2018 gautamUpdateALSFibers switched out

A follow-up on the discussion from today's lunch meeting - Rana pointed out that rotation of the fiber in the mount by ~5degrees cannot account for such large power fluctuations. Here is a 3 day trend from my polarization monitoring setup. Assuming the output fiber coupler rotates in its mount by 5 degrees, and assuming the input light is aligned to one of the fiber's special axes, then we expect <1% fluctuation in the power. But the attached trend shows much more drastic variations, more like 25% in the p-polarization (which is what I assume we use for the beat, since the majority of light is in this polarization, both for PSL and EX). I want to say that the periodicity in the power fluctuations is ~12hours, and so this fluctuation is somehow being modulated by the lab temperature, but unfortunately, we don't have the PSL enclosure temperature logged in order to compare coherence.

Steve: your  plots look like temperature driven


The "beat length" of the fiber is quoted as <=2.7mm. This means that a linearly polarized beam that is not oriented along one of the special axes of the fiber will be rotated through 180 degrees over 2.7mm of propagation through the fiber. I can't find a number for the coefficient of thermal expansion of the fiber, but if temperature driven fluctuations are changing the length of the fiber by 300um, it would account for ~12% power fluctuation between the two polarizations in the monitoring setup, which is in the ballpark we are seeing...

  13756   Tue Apr 17 09:57:09 2018 SteveUpdateGeneralseismometer interfaces

 

Quote:

I've been looking into recovering the seismic BLRMs for the BS Trillium seismometer. It looks like the problem is probably in the anti-aliasing board. There's some heavy stuff sitting on top of it in the rack, so I'll take a look at it later when someone can give me a hand getting it out.

In detail, after verifying that there are signals coming directly out of the seismometer, I tried to inject a signal into the AA board and see it appear in one of the seismometer channels.

  1. I looked specifically at C1:PEM-SEIS_BS_Z_IN1 (Ch9), C1:PEM-SEIS_BS_X_IN1 (Ch7), and C1:PEM-ACC_MC2_Y_IN1 (Ch27). All of these channels have between 2000--3000 cts.
  2. I tried injecting a 200 mVpp signal at 1.7862 Hz into each of these channels, but the the output did not change.
  3. All channels have 0 cts when the power to the AA board is off.
  4. I then tried to inject the same signal into the AA board and see it at the output. The setup is shown in the first attachment. The second BNC coming out of the function generator is going to one of the AA board inputs; the 32 pin cable is coming directly from the output. All channels give 4.6 V when when the board is powered on regardless of wheter any signal is being injected.
  5. To verify that the AA board is likely the culprit, I also injected the same signals directly into the ADC. The setup is shown in the second attachment. The 32 pin cable is going directly to the ADC. When injecting the same signals into the appropriate channels the above channels show between 200--300 cts, and 0 cts when no signal is injected.

 

  13755   Mon Apr 16 22:09:53 2018 KevinUpdateGeneralpower outage - BLRM recovery

I've been looking into recovering the seismic BLRMs for the BS Trillium seismometer. It looks like the problem is probably in the anti-aliasing board. There's some heavy stuff sitting on top of it in the rack, so I'll take a look at it later when someone can give me a hand getting it out.

In detail, after verifying that there are signals coming directly out of the seismometer, I tried to inject a signal into the AA board and see it appear in one of the seismometer channels.

  1. I looked specifically at C1:PEM-SEIS_BS_Z_IN1 (Ch9), C1:PEM-SEIS_BS_X_IN1 (Ch7), and C1:PEM-ACC_MC2_Y_IN1 (Ch27). All of these channels have between 2000--3000 cts.
  2. I tried injecting a 200 mVpp signal at 1.7862 Hz into each of these channels, but the the output did not change.
  3. All channels have 0 cts when the power to the AA board is off.
  4. I then tried to inject the same signal into the AA board and see it at the output. The setup is shown in the first attachment. The second BNC coming out of the function generator is going to one of the AA board inputs; the 32 pin cable is coming directly from the output. All channels give 4.6 V when when the board is powered on regardless of wheter any signal is being injected.
  5. To verify that the AA board is likely the culprit, I also injected the same signals directly into the ADC. The setup is shown in the second attachment. The 32 pin cable is going directly to the ADC. When injecting the same signals into the appropriate channels the above channels show between 200--300 cts, and 0 cts when no signal is injected.
  13754   Sat Apr 14 14:42:09 2018 gautamUpdateALSFibers switched out

It looks like the drift in polarization content in the PSL pickoff is actually much higher than that in the EX pickoff. Note that to prevent the P-pol diode from saturating, I put an ND filter in front of the PD, so the Y axis actually has to be multiplied by 10 to compare power in S and P polarizations. If this drift is because of the input (linear) polarization being poorly matched to one of the fiber's special axes, then we can improve the situation relatively easily. But if the polarization drift is happening as a result of time-varying stress (due to temp. fluctuations, acoustics etc) on the (PM) fiber from the PSL fiber coupler to the BeatMouth, then I think this is a much more challenging problem to solve.

I'll attempt to quantify the contribution (in Hz/rtHz) of beat amplitude RIN to phase tracker output noise, which will tell us how much of a problem this really is and in which frequency bands. In particular, I'm interested in seeing if the excess noise around 100 Hz is because of beat amplitude fluctuations. But on the evidence thus far, I've seen the beat amplitude drift by ~15 dB (over long timescales) on the control room network analyzer, and this drift seems to be dominated by PSL light amplitude fluctuations.

  13753   Fri Apr 13 17:56:26 2018 gautamUpdateALSFibers switched out

I swapped the EX fiber for the PSL fiber in the polarization monitoring setup. There is a lot more power in this fiber, and one of the PDs was saturated. I should really have put a PBS to cut the power, but I opted for putting an absorptive ND1.0 filter on the PD instead for this test. I want to monitor the stability in this beam and compare it to the EX beam's polarization wandering.

  13752   Fri Apr 13 16:59:12 2018 gautamUpdateALSEX green mode-matching

THIS CALCULATION IS WRONG FOR THE OVERCOUPLED CAV.

Summary:

Mode-matching efficiency of EX green light into the arm cavity is ~70*%, as measured using the visibility. 

Details:

I wanted to get an estimate for the mode-matching of the EX green beam into the arm cavity. I did the following:

  1. Locked arm cavities to IR. Ran dither alignment servos to maximize the transmission of IR on both arms. The X arm dither alignment servo needs some touching up, I can achieve higher TRX by hand than by running the dither.
  2. Aligned green PZT mirrors so as to maximize GTRX. Achieved level as 0.47.
  3. Went to EX table and tweaked the two available mode-matching lens positions on their translational stages. Saw a quadratic maximum of GTRX about some equilibrium position (where the lenses are now).
  4. Measured average value of the green PDH reflection DC level whiel green TEM00 mode was locked. P_{\mathrm{locked}} = 716 \mathrm{cts}.
  5. Misaligned ITMX macroscopically. Measured the average value of the green PDH reflection DC level again. P_{\mathrm{misaligned}} = 3800 \mathrm{cts}.
  6. Closed EX Green shutter. Measured the average value of the green PDH reflection DC level. P_{\mathrm{dark}} = 30 \mathrm{cts}.
  7. Modulation depth of the EX PDH was determined to be 90mrad. Based on this, power in sideband is negligible compared to power in the carrier, so I didn't bother correcting for sideband power in reflection.
  8. Mode-matching efficiency calculated as \frac{P_{\mathrm{misaligned}} - P_{\mathrm{locked}}}{P_{\mathrm{misaligned}} + P_{\mathrm{locked}} - 2P_{\mathrm{dark}} }.

Comments:

This amount of mode-matching is rather disappointing - using a la mode, the calculated mode-matching efficiency is nearly 100%, but 70% is a far cry from this. The fact that I can't improve this number by either tweaking the steering or by moving the MM lenses around suggests that the estimate of the target arm mode is probably incorrect (the non-gaussianity of the input beam itself is not quantified yet, but I don't believe this input beam can account for 30% mismatch). For the Y-arm, the green REFL DC level is actually higher when locked than when ITMY is misaligned. WTF?? surpriseOnly explanation I can think of is that the PD is saturated when green is unlocked - but why does the ADC saturate at ~3000cts and not 32000?


This data is almost certainly bogus as the AA box at 1X9 is powered by +/-5VDC and not +/-15VDC. I didn't check but I believe the situation is the same at the Y-end.

3000 cts is ~1V into the ADC. I am going to change the supply voltage to this box (which also reads in ETMX OSEMS) to +/-15V so that we can use the full range of the ADC.


gautam Apr 26 630pm: I re-did the measurement by directly monitoring the REFLDC on a scope, and the situation is not much better. I calculate a MM of 70% into the arm. This is sensitive to the lens positions - while I was working on the EX fiber coupling, I had bumped the lens mounted on a translational stage on the EX table lightly, and I had to move that lens around today in order to recover the GTRX level of 0.5 that I am used to seeing (with arm aligned to maximize IR transmission). So there is definitely room for optimization here.


 

  13751   Fri Apr 13 11:02:41 2018 gautamUpdateALSEX fiber polarization drift

Attachment #1 shows the drift of the polarization content of the light from EX entering the BeatMouth. Seems rather large (~10%). I'm going to tweak the X end fiber coupling setup a bit to see if this can be improved. This performance is also a good benchmark to compare the PSL IR light polarization drift. I am going to ask Steve to order Thorlabs K6XS, which has a locking screw for the rotational DoF. With this feature, and by installing some HWPs at the input coupling point, we can ensure that we are coupling light into one of the special axes in a much more deterministic way. 

  13750   Fri Apr 13 00:20:46 2018 ranaUpdatePEMMEDM setup

changed the setpoint of the EX Seismomter T ctrl servo from 35 to 39 C to see if this helps the stability by decreasing the cooldown time constant.

  13749   Thu Apr 12 18:12:49 2018 gautamUpdateALSNPRO channels hijacked

Summary:

  1. Today, the measured IR ALS noise for the X arm was dramatically improved. The main change was that I improved the alignment of the PSL pickoff beam into its fiber coupler.
  2. The noise level was non-stationary, leading me to suspect power modulation of the RF beat amplitude.
  3. I am now measuring the stability of the power in the two polarizations coming from EX table to the PSL table, using the PSL diagnostic connector channels.
  4. The EX beam is S-polarized when it is coupled into the fiber. The PSL beam is P-polarized. However, it looks like I have coupled light along orthogonal axes into the fiber, such that when the EX light gets to the PSL table, most of it is in the P-polarization, as judged by my PER measurement setup (i.e. the alignment keys at the PSL table and at the EX table are orthogonal). So it still seems like there is something to be gained by trying to improve the PER a bit more.

Details:

Today, I decided to check the power coupled into the PSL fiber for the BeatMouth. Surprisingly, it was only 200uW, while I had ~3.15mW going into it in January. Presumably some alignment drifting happened. So I re-aligned the beam into the fiber using the steering mirror immediately before the fiber coupler. I managed to get ~2.9mW in without much effort, and I figured this is sufficient for a first pass, so I didn't try too much more. I then tried making an ALS beat spectrum measurement (arm locked to IMC length using POX, green following the arm using end PDH servo). Surprisingly, the noise performannce was almost as good as the reference! See Attachment #1, in which the red curve is an IR beat (while all others are green beats). The Y arm green beat performance isn't stellar, but one problem at a time. Moreover, the kind of coherence structure between the arm error signal and the ALS beat signal that I reported here was totally absent today.

Upon further investigation, I found that the noise level was actually breathing quite significantly on timescales of minutes. While I was able to successfully keep the TEM00 mode of the PSL beam resonant inside the arm cavity by using the ALS beat frequency as an error signal and MC2 as a frequency actuator, the DC stability was very poor and TRX was wandering around by 50%. So my new hypothesis is that the excess ALS noise is because of one or more of

  • Beam jitter at coupling point into fiber.
  • Polarization drift of the IR beams.

While I did some work in trying to align the PSL IR pickoff into the fiber along the fast (P-pol) axis, I haven't done anything for the X end pickoff beam. So perhaps the fluctuations in the EX IR power is causing beatnote amplitude fluctuations. In the delay line + phase tracker frequency discriminator, I think RF beatnote amplitude fluctuations can couple into phase noise linearly. For such an apparently important noise source, I can't remeber ever including it in any of the ALS noise budgets.

Before Ph237 today, I decided to use my polarization monitoring setup and check the "RIN" of power in the two polarizations coming out of the fiber on the PSL table. For this purpose, I decided to hijack the Acromag channels used for the PSL diagnostics connector Attachment #2 shows that there is fluctuations at the level of ~10% in the p-polarization. This is the "desired" polarization in that I aligned the PSL beam into the fiber to maximize the power in this polarization. So assuming the power fluctuations in the PSL beam are negligible, this translates to sqrt(10) ~3% fluctuation in the RF beat amplitude. This is at best a conservative estimate, as in reality, there is probably more AM because of the non PM fibers inside the beatmouth.

All of this still doesn't explain the coherence between the measured ALS noise and the arm error signal at 100s of Hz (which presumably can only happen via frequency noise in the PSL).

Another "mystery" - yesterday, while I was working on recovering the Y arm green beat signal on the PSL table, I eventually got a beat signal that was ~20mVpp into 50ohms, which is approximately the same as what I measured when the Y arm ALS performance was "nominal", more than a year ago. But while viewing the Y arm beats (green and IR) simultaneously on an o'scope, I wasn't able to keep both signals synchronised while triggering on one (even though the IR beat frequency was half the green beat frequency). This means there is a huge amount of relative phase noise between the green and IR beats. What (if anything) does this mean? The differential noise between these two beats should be (i) phase noise at the fiber coupler/ inside the fiber itself, and (ii) scatter noise in the green light transmitted through the cavity. Is it "expected" that the relative phase noise between these two signals is so large that we can't view both of them on a common trigger signal on an o'scope? surpriseAlso - the green mode-matching into the Y arm is abysmal.

Anyways - I'm going to try and tweak the PER and mode-matching into the X end fiber a little and monitor the polarization stability (nothing too invasive for now, eventually, I want to install the new fiber couplers I acquired but for now I'll only change alignment into and rotation of the fiber coupler on the EX table). It would also be interesting to compare my "optimized" PSL drift to the unoptimized EX power drift. So the PSL diagnostic channels will not show any actual PSL diagnostic information until I plug it back in. But I suspect that the EPICS record names and physical channel wiring are wrong anyways - I hooked up my two photodiode signals into what I would believe is the "Diode 1 Power" and "Laser crystal temperature" monitors (as per the schematic), but the signals actually show up for me in "Diode 2 Power" (p-pol) and "Didoe 1 Temperature" (s-pol).

Annoyingly, there is no wiring diagram - on my todo list i guess...

@Steve - could you please take a photo of the EX table and update the wiki? I think the photo we have is a bit dated, the fiber coupler and transmon PDs aren't in it...

  13748   Thu Apr 12 10:15:33 2018 KiraUpdatePEMMEDM setup

Another update. I've changed the on/off button so that it's visible which state it's in. I did that by changing the type of C1:PEM-SEIS-EX_TEMP_SLOWLOOP from ai to bi (I checked the FSS script and copied the entry for the slowloop). Previously, MEDM was giving me an error that it wasn't an ENUM value when I wanted to use a choice button to indicate the value of slowloop, and this solved the issue. I've also added a StripTool button.

  13747   Wed Apr 11 10:47:26 2018 SteveUpdateSUSsatellite amps labeled

Satellite amplifiers labeled with date. Old labels left on.

  13746   Wed Apr 11 01:34:31 2018 gautamUpdateIOOActivities today

[kevin, gautam]

activities done today - details/plots/data/evidence tomorrow.

  1. Checked XARM loop shape. Updated NB code to fetch POX data from NDS and undo loop shape rather than using calibration filter bank.
  2. Checked POX loop calibration (m/ct). Number I was using was 8e-13. Tonight we measured 9e-13. Updated filter bank.
  3. Tried to get Y arm green ALS going.
    • Improved GTRY from ~0.05 to 0.3.
    • Tried to improve mode matching onto BBPD on PSL table to see a green beat.
    • But we were unsuccessful.
    • I think I got the near and far field alignment right, and the EY laser temp is set such that I can see an IR beat @~30MHz (so green beat should be at 60 MHz).
    • But I couldn't see anything with scope or with HP spec analyzer.
    • More tomorrow. Motivation to get green ALS working is to get some more confidence that the excess noise is indeed on the PSL light.
  13745   Tue Apr 10 15:42:08 2018 KiraUpdatePEMMEDM setup

An update to the screen. I changed the min/max values for some of the parameters, as well as changing the script so that I could specify the integral gain in terms of 1e-5. I've also added this screen to the PEM tab in the sitemap.

  13744   Tue Apr 10 14:28:44 2018 gautamUpdateIOOFurther debugging

I am working on IMC electronics. IMC is misaligned until further notice.

  13743   Mon Apr 9 23:49:50 2018 ranaUpdatePSLPSL chans

I think we can scrap the 126MOPA channels since they're associated with the Lightwave NPRO and MOPA. We should add the channels that we need for monitoring the Innolight NPRO from the d-sub connector on its controller.

  13742   Mon Apr 9 23:28:49 2018 johannesConfigurationDAQc1psl channel list

I made a list of all the physical c1psl channels to get a better idea for how many acromags we need to replace it eventually. There  3123 unit is the one whose failure had prevented c1psl from booting, which is why it was unplugged (elog post 12852), and its channels have been inactive since. Are the 126MOPA channels used for the current mephisto? 126 tells me it's for an old lightwave laser, but I was checking a few and found that they have non-zero, changing values, so they may have been rewired.

It also hosts some virtual channels for the ISS with root C1:PSL-ISS_ defined in iss.db and dc.db, the PSL particle counter with root C1:PEM- defined in PCount.db  and a whole lot of PSL status channels defined in pslstatus.db. Transfering these virtual channels to a different machine is almost trivial, but the serial readout of the particle counter would have to find a new home.

Long story short - we need:

Function Type # Channels #Channels (no MOPA) # Units # Units (no MOPA)
ADC XT1221 34 21 5 3
DAC XT1541 17 14 3 2
BIO XT1111 19 10 2 1

 



3113 - ADC

C1:PSL-126MOPA_126PWR
C1:PSL-126MOPA_DTMP
C1:PSL-126MOPA_LTMP
C1:PSL-126MOPA_DMON
C1:PSL-126MOPA_LMON
C1:PSL-126MOPA_CURMON
C1:PSL-126MOPA_DTEC
C1:PSL-126MOPA_LTEC
C1:PSL-126MOPA_CURMON2
C1:PSL-126MOPA_HTEMP
C1:PSL-126MOPA_HTEMPSET
C1:PSL-FSS_RFPDDC
C1:PSL-FSS_LODET
C1:PSL-FSS_FAST
C1:PSL-FSS_PCDRIVE
C1:PSL-FSS_MODET
C1:PSL-FSS_VCODETPWR
C1:PSL-FSS_TIDALOUT
C1:PSL-PMC_RFPDDC
C1:PSL-PMC_LODET
C1:PSL-PMC_PZT
C1:PSL-PMC_MODET


3123 - ADC (failed)

C1:PSL-126MOPA_AMPMON
C1:PSL-126MOPA_126MON
C1:PSL-FSS_RCTRANSPD
C1:PSL-FSS_MINCOMEAS
C1:PSL-FSS_RMTEMP
C1:PSL-FSS_RCTEMP
C1:PSL-FSS_MIXERM
C1:PSL-FSS_SLOWM
C1:PSL-FSS_TIDALINPUT
C1:PSL-PMC_PMCTRANSPD
C1:PSL-PMC_PMCERR
C1:PSL-PPKTP_TEMP


4116 - DAC

C1:PSL-126MOPA_126CURADJ
C1:PSL-126MOPA_DCAMP
C1:PSL-126MOPA_DCAMP-
C1:PSL-FSS_INOFFSET
C1:PSL-FSS_MGAIN
C1:PSL-FSS_FASTGAIN
C1:PSL-FSS_PHCON
C1:PSL-FSS_RFADJ
C1:PSL-FSS_SLOWDC
C1:PSL-FSS_VCOMODLEVEL
C1:PSL-FSS_TIDAL
C1:PSL-FSS_TIDALSET
C1:PSL-PMC_GAIN
C1:PSL-PMC_INOFFSET
C1:PSL-PMC_PHCON
C1:PSL-PMC_RFADJ
C1:PSL-PMC_RAMP


XVME-210 - Binary Input

C1:PSL-126MOPA_FAULT
C1:PSL-126MOPA_INTERLOCK
C1:PSL-126MOPA_SHUTTER
C1:PSL-126MOPA_126LASE
C1:PSL-126MOPA_AMPON


XVME-220 - Binary Output

C1:PSL-126MOPA_126NE
C1:PSL-126MOPA_126STANDBY
C1:PSL-126MOPA_SHUTOPENEX
C1:PSL-126MOPA_STANDBY
C1:PSL-FSS_SW1
C1:PSL-FSS_SW2
C1:PSL-FSS_FASTSWEEP
C1:PSL-FSS_PHFLIP
C1:PSL-FSS_VCOTESTSW
C1:PSL-FSS_VCOWIDESW
C1:PSL-PMC_SW1
C1:PSL-PMC_SW2
C1:PSL-PMC_PHFLIP
C1:PSL-PMC_BLANK

  13741   Mon Apr 9 18:46:03 2018 gautamUpdateIOOFurther debugging
  1. I analyzed the data from the free swinging MC test conducted over the weekend. Attachment #1 shows the spectra. Color scheme is same for all panels.
    • I am suspicious of MC3: why does the LR coil see almost no Yaw motion?
    • The "equilibrium" values of all the sensor signals (at the IN1 of the coil input filters) are within 20% of each other (for MC3, but also MC1 and MC2).
    • The position resonance is also sensed more by the side coil than by the LR coil.
    • To rule out satellite box shenanigans, I just switched the SRM and MC3 satellite boxes. But coherence between frequency noise as sensed by the arms remain.
  2. I decided to clean up my IMC nosie budget a bit more.
    • Attachment #2 shows the NB as of today. I'll choose a better color palette for the next update.
    • "Seismic" trace is estimated using the 40m gwinc file - the MC2 stack is probably different from the others and so it's contribution is probably more, but I think this will suffice for a first estimate.
    • "RAM" trace is measured at the CM board input, with MC2 misaligned.
    • The unaccounted noise is evident from above ~8 Hz.
    • More noises will be added as they are measured.
    • I am going to spend some time working on modeling the CM board noise and TF in LTspice. I tried getting a measurement of the transfer function fron IN1 to the FAST output of the CM board with the SR785 (motivation being to add the contribution of the input referred CM board noise to the NB plot), but I suspect I screwed up something w.r.t. the excitation amplitude, as I am getting a totally nonsensical shape, which also seems to depend on my input excitation amplitude. I don't think the output is saturated (viewed during measurement on a scope), but perhaps there are some subtle effects going on.
  13740   Mon Apr 9 16:30:21 2018 KiraUpdatePEMMEDM setup

I created an MEDM screen for the PID control. In addition, I added a new EPICS channel for the setpoint so that it could be adjusted using the MEDM screen.

Edit: forgot to mention the channel name is C1:PEM-SEIS_EX_TEMP_SETPOINT

Edit #2: the path for the MEDM is /opt/rtcds/caltech/c1/medm/c1pem/C1PEM_SEIS_EX_TCTRL.adl

  13739   Mon Apr 9 08:39:39 2018 SteveUpdatePEMM5.3 eq Souther CA

Earth quake M5.3    2018-04-05 19:29:16UTC          Santa Cruz Island, CA

 

  13738   Fri Apr 6 22:23:53 2018 KevinUpdateIOOCoil driver noise

 

Quote:

Why is MC2 LR so different from the others???

The previous measurements were made from the coil driver output monitors. To investigate why the MC2 LR coil has less noise than the other coils, I also measured the noise at the output to the coils.

The circuit diagram for the coil driver board is given in D010001 and a picture of the rack is on the 40m wiki here. The coil driver boards are in the upper left quadrant of the rack. The input to the board is the column of LEMOs which are the "Coil Test In" inputs on the schematic. The output monitors are the row of LEMOs to the right of the input LEMOs and are the "FP Coil Volt Mon" outputs on the schematic. The output to the coils "Coil Out" in the schematic are carried through a DB15 connector.

The attachment shows the voltage noise for the MC2 LR coil as well as the UL coil which is similar to all of the other coils measured in the previous measurement. While the LR coil is less noisy than the UL coil as measured at the output monitor, they have the same noise spectrum as measured at the output to the coils themselves. So there must be something wrong with the buffer circuit for the MC2 LR voltage monitor, but the output to the coils themselves is the same as for the other coils.

  13737   Fri Apr 6 21:39:09 2018 gautamUpdateIOOMC2 suspension health checkup

While Kevin is working on the MC2 electronics chain - we disconnected the output to the optic (DB15 connector on coil driver board). I decided to look at the 'free' freeswinging MC2 OSEM shadow sensor data. Attachment #1 suggests that the suspension eigenmodes are showing up in the shadow sensors, but the 0.8Hz peak seems rather small, especially compared to the result presented in this elog.

Maybe I'll kick all 3 MC optics tonight and let them ringdown overnight, may not be a bad idea to checkup on the health of the MC suspensions/satellite boxes... [MC suspensions were kicked @1207113574]. PSL shutter will remain closed overnight...

Quote:

Why is MC2 LR so different from the others???

 

  13736   Fri Apr 6 18:28:57 2018 ranaUpdatePEMPID tuning

Made some changes:

  1. Set P and D gains to zero. We only need slow drift control.
  2. Changed names of the python script and .ini file to distinguish it from the FSS stuff. Lives in scripts/PEM/
  3. removed debug flag from argParse. To run in non-debug mode you use the "-O" option of python as usual.
  4. Fixed the upper/lower limit convention for the heater. Was backwards.
  5. Removed the "rail" function that was defined. We can just use numpy.clip since that's already built in.

There is also now a StripTool file in the scripts/PEM directory which has appropriate channel names and scales for PID loop tuning. Use this file!

I'm leaving it running over the weekend with K_I = -0.003. There is a StripTool on rossa which you can watch. The code itself is running on a tmux session on megatron. Let's ONLY run this code there until we're satisfied that things are good.

Update Sun Apr 8 00:40:11 2018: Lowered gain by factors of 3 down to -0.0001 Saturday afternoon. Seems like still oscillating a bit, now with a ~4 hour period. Setting it to -3e-5 now. Usually we have a linear feedback loop, but our actuation voltage actually gets squared (P = I^2 R) before being integrated to produce temperature. Wonder if we should think of linearizing the feedback control signal to make the loop act nicer.

Update Sun Apr 8 21:09:48 2018: Set K_I = -1e-5 earlier today. Seems to have stabilized nearly, but temperature swings are still +/- 1 K. Will need to add some proportional feedback (K_P) to increase the loop bandwidth, but system is at least sort of stable now. Probably should start construction of EY,BS systems now.

  13735   Fri Apr 6 16:17:20 2018 KiraUpdatePEMPID tuning

I have been trying to tune the PID and have managed to descrease the oscillations without saturating the actuator. I'm going to model the system to calculate the exact values of P, I and D in order to get rid of the oscillations altogether. I was going to record the data using Data Viewer, but there seems to be some issue with that, so I'm using StripTool for now.

  13734   Fri Apr 6 16:07:18 2018 gautamUpdateCDSFrequent EPICS dropouts

Kira informed me that she was having trouble accessing past data for her PID tuning tests. Looking at the last day of data, it looks like there are frequent EPICS data dropouts, each up to a few hours. Observations (see Attachment #1 for evidence):

  1. Problem seems to be with writing these EPICS channels to frames, as the StripTool traces do not show the same discontinuities.
  2. Problem does not seem local to c1auxex (new Acromag machine). These dropouts are also happening in other EPICS channel records. I have plotted PMC REFL, which is logged by the slow machine c1psl, and you can see the dropouts happen at the same times.
  3. Problem does not seem to happen to the fast channel DQ records (see ETMX Sensor record plotted for 24 hours, there are no discontinuties).

It is difficult to diagnose how long this has been going on for, as once you start pulling longer stretches of data on dataviewer, any "data freezes" are washed out in the extent of data plotted.

  13733   Fri Apr 6 10:00:29 2018 gautamUpdateCDSCDS puzzle
Quote:

Clearly, there is a discrepancy for f>20Hz. Why?

Spectral leakage

  13732   Thu Apr 5 19:31:17 2018 gautamUpdateCDSEPICS processes for c1ioo dead

I found all the EPICS channels for the model c1ioo on the FE c1ioo to be blank just now. The realtime model itself seemed to be running fine, judging by the IMC alignment (as the WFS loops seemed to still be running okay). I couldn't find any useful info in demsg but I don't know what I'm looking for. So my guess is that somehow the EPICS process for that model died. Unclear why.

  13731   Thu Apr 5 13:46:42 2018 gautamUpdateSUSbig earthquake

Seems like there was a 5.3 magnitude EQ ~10km from us (though I didn't feel it). All watchdogs were tripped so our mirrors definitely felt it. ITMX is stuck (but all the other optics are damping fine). I tried the usual jiggling of DC bias voltage but ITMX still seems stuck. Probably a good sign that the magnet hasn't come off, but not ideal that I can't shake it free..

edit: after a bit more vigorous shaking, ITMX was freed. I had to move the bias slider by +/-10,000 cts, whereas initially I was trying +/-2000 cts. There is a tendency for the optic to get stuck again once it has been freed (while the optic's free swinging motion damps out), so I had to keep an eye out and as soon as the optic was freed, I re-engaged the damping servos to damp out the optic motion quickly.

  13730   Thu Apr 5 12:13:18 2018 KojiUpdateIOOCoil driver noise

Why is MC2 LR so different from the others???

  13729   Thu Apr 5 10:38:38 2018 gautamUpdateCDSCDS puzzle

I'm probably doing something stupid - but I've not been able to figure this out. In the MC1 and MC3 coil driver filter banks, we have a digital "28HzELP" filter module in FM9. Attachment #1 shows the MC1 filterbanks. In the shown configuration, I would expect the only difference between the "IN1" and "OUT" testpoints to be the transfer function of said ELP filter, after all, it is just a bunch of multiplications by filter coefficients. But yesterday looking at some DTT traces, their shapes looked suspicious. So today, I did the analysis entirely offline (motivation being to rule out DTT weirdness) using scipy's welch. Attachment #2 shows the ASDs of the IN1 and OUT testpoint data (collected for 30s, fft length is set to 2 seconds, and hanning window from scipy is used). I've also plotted the "expected" spectral shape, by loading the sos coefficients from the filter file and using scipy to compute the transfer function.

Clearly, there is a discrepancy for f>20Hz. Why?

Code used to generate this plot (and also a datafile to facilitate offline plotting) is attached in the tarball Attachment #3. Note that I am using a function from my Noise Budget repo to read in the Foton filter file...

*ChrisW suggested ruling out spectral leakage. I re-ran the script with (i) 180 seconds of data (ii) fft length of 15 seconds and (iii) blackman-harris window instead of Hanning. Attachment #4 shows similar discrepancy between expectation and measurement...

  13728   Thu Apr 5 04:36:56 2018 KevinUpdateIOOCoil driver noise

[Gautam, Kevin]

We measured the MC coil driver noise at the output monitors of the coil driver board with an SR785 in order to further diagnose the excess IMC frequency noise.

Attachments 1 and 2 show the noise for the UL coils of MC3 and MC2 with various combinations of output filters engaged. When the 28 Hz elliptic filter is on, the analog dewhitening filter is off, and vice versa. The effect of the analog low pass filter is visible in MC3, but the effect of the digital low pass filter is swamped by the DAC noise.

We locked the arms and measured the ALS beatnote in each of these filter combinations, but which filters were on did not effect the excess IMC frequency noise. This suggests that the coil drivers are not responsible for the excess noise.

Attachment 2 shows the noise for all five coils on MC1, MC2, and MC3 as well as for ITMY, which is on a different DAC card from the MCs. All filters were on for these measurements.

 

  13727   Wed Apr 4 16:23:39 2018 gautamUpdateCDSslow machine bootfest

[johannes, gautam]

It's been a while - but today, all slow machines (with the exception of c1auxex) were un-telnetable. c1psl, c1iool0, c1susaux, c1iscaux1, c1iscaux2, c1aux and c1auxey were rebooted. Usual satellite box unplugging was done to avoid ITMX getting stuck.

  13726   Wed Apr 4 16:23:10 2018 KiraUpdatePEMPID test

I did a step response for the loop from 35 degrees to 40 degrees. The PID is not properly tuned, so the signal oscillates. In the graph, the blue curve is the temperature of the can in celcius and the green curve is the heating power in watts. The x-axis is in minutes. Before, the signal was too noisy to do a proper step response, so I placed a 3.3 microF capacitor in parallel with the resistor in my temperature sensor circuit (I'll draw and attach this updated version). This created a 5 Hz low pass filter and the signal is now pretty clean.

-----

I also added in new Epics channels so that we could log the data using Data Viewer. The channels I added were C1:PEM-SEIS_EX_TEMP_MON_CELCIUS and C1:PEM-SEIS_EX_TEMP_CTRL_WATTS. I used 13023 as a guide on how to do this.

Update: the channels work and show data in Data Viewer

-----

Edit: I've attached a photo of the circuit with the capacitor indicated. It is in parallel with the resistor below it. I've attached an updated circuit diagram as well.

  13725   Mon Apr 2 15:14:21 2018 KojiUpdateGeneralModulation depth measurement for an aLIGO EOM

The new matching circuit was tested.

Results:

f_nominal  f_actual  response    required mod.  drivng power
 [MHz]      [MHz]    [mrad/V]     [rad]         needed [dBm]
  9.1       9.1        55         0.22      =>   22
118.3     118.2        16         0.01      =>    6

 45.5      45.4        45         0.28      =>   25
 24.1       N/A         2.1       0.014     =>   27

Comments:

- 9.1MHz and 118.3MHz: They are just fine.

- 24.1MHz: Definitely better (>x3) than the previous trial to combine 118MHz & 24MHz.
  We got about the same modulation with the 50Ohm terminated bare crystal (for the port1).
  So, this is sort of the best we can do for the 24.1MHz with the current approach.
  The driving power of 27dBm is required at 24.1MHz

- About the 45MHz
  - The driving power of 27dBm is required at 24.1MHz
  - The maximum driving power with the AM stabilized driver is 23dBm, nominally to say.
  - I wonder how we can reduce resistance (and capacitance) of the 45MHz further...?
  - I also wonder if the IFO can be locked with reduced modulation (0.28 rad->0.2 rad)
  - Can the driver max power be boosted a bit? (i.e. adding an attenuator in the RF power detection path)

 

  13724   Fri Mar 30 22:37:36 2018 KevinUpdateIOOMCREFL_PD Optical response measurement

[Gautam, Kevin]

We redid the measurement measuring the voltage noise from the REFL PD demod board output monitor with an SR785 with the noise eater on and off. We used a 100x preamp to amplify the signal above the SR785 noise. The SR785 noise floor was measured with the input to the preamp terminated with 50 ohms. The spectra shown have been corrected for the 100x amplification.

This measurement shows no difference with the noise eater on or off.

Quote:

the noise eater on/off measurements should be done for 0-100 kHz and from the demod board output monitor

 

  13723   Fri Mar 30 16:10:46 2018 KiraUpdatePEMPID test

I created two new channels today, C1:PEM-SEIS_EX_TEMP_MON_CELCIUS, which turns the output voltage signal into degrees C, and C1:PEM-SEIS_EX_TEMP_CTRL_WATTS, which takes the input voltage from the DAC and turns it into a value of watts. I'm trying to stabilize the temperature at 35 degrees, but it's taking a lot longer than expected. Perhaps we'll need to use different values for P and I and decrease the noise in the sensor, since right now there's about a 10 degree variation between the highest and lowest values.

  13722   Fri Mar 30 06:16:45 2018 ranaUpdatePEMPID test

Can't really figure out what this plot means. We need to see the sensor (in units of deg C) and the control signal (in heating power (W)). The plot should show a few step responses with the PID loop on, so that we can see the loop response time. Please zoom in on the axes so that we can see what's happening.

Quote:

[Kira, Gautam]

We closed the loop today and implemented the PID script. I have attached the StripTool graph for an integral value of 0.5 and proportional value of 20. We had some issues getting it to work properly and it would oscillate between some low values of the control voltage. The set point here was -3.20, which corresponds to about a 20 degree increase in temperature. The next step would be to find which values of Kp, Ki, and Kd would work in this case and low pass filter the signal from the temperature sensor, and also create an MEDM screen for easier PID control.

 

  13721   Fri Mar 30 06:14:31 2018 ranaUpdateIOOMCREFL_PD Optical response measurement

the noise eater on/off measurements should be done for 0-100 kHz and from the demod board output monitor

  13720   Fri Mar 30 03:23:50 2018 KojiUpdateGeneralaLIGO EOM work

I have been working on the aux beat setup on the PSL table between 9PM-3AM.

This work involved:

- Turning off the main marconi
- Turning off the freq generation unit (incl IMC modulation)
- Closing the PSL shutter

After the work, these were reverted and the IMC and both arms have been locked.

  13719   Thu Mar 29 17:57:36 2018 arijitUpdateIOOMCREFL_PD Optical response measurement

Kevin, Gautam and Arijit

Today we performed the in-loop noise measurements of the MCREFL-PD using the SR785 to ascertain the effect of the Noise Eater on the laser. We took the measurements at the demodulated output channel from the MCREFL-PD. We performed a series of 6 measurements with the Noise Eater ''ON'' and ''OFF''. The first data set is an outlier probably, due to some transient effects. The remaining data sets were recorded in succession with a time interval of 5 minutes each between the Noise Eater in the ''ON'' and ''OFF'' state. We used the calibration factor of 13kHz/Vrms from elog 13696 to convert the V_rms to Hz-scale.

The conclusion is that the NOISE EATER does not have any noticeable effect on the noise measurements.

ALS beat spectrum and also the arm control signal look as they did before. coherence between arm control signals (in POX/POY lock) is high between 10-100Hz, so looks like there is still excess frequency noise in the MC transmitted light. Looking at POX as an OOL sensor with the arm under ALS control shows ~10x the noise at 100 Hz compared to the "nominal" level, consistent with what Koji and I observed ~3weeks ago.

We tried swapping out Marconis. Problem persists. So Marconi is not to blame. I wanted to rule this out as in Jan, Steve and I had installed a 10MHz reference to the rear of the Marconi.

  13718   Thu Mar 29 17:14:42 2018 KiraUpdatePEMPID test

[Kira, Gautam]

We closed the loop today and implemented the PID script. I have attached the StripTool graph for an integral value of 0.5 and proportional value of 20. We had some issues getting it to work properly and it would oscillate between some low values of the control voltage. The set point here was -3.20, which corresponds to about a 20 degree increase in temperature. The next step would be to find which values of Kp, Ki, and Kd would work in this case and low pass filter the signal from the temperature sensor, and also create an MEDM screen for easier PID control.

  13717   Thu Mar 29 12:03:37 2018 Jon RichardsonSummaryGeneralProof-of-Concept SRC Gouy Phase Measurement

I've been developing an idea for making a direct measurement of the SRC Gouy phase at RF. It's a very different approach from what has been tried before. Prior to attempting this at the sites, I'm interested in making a proof-of-concept measurement demonstrating the technique on the 40m. The finesse of the 40m SRC will be slightly higher than at the sites due to its lower-transmission SRM. Thus if this technique does not work at the 40m, it almost certainly will not work at the sites.

The idea is, with the IFO locked in a signal-recycled Michelson configuration (PRM and both ETMs misaligned), to inject an auxiliary laser from the AS port and measure its reflection from the SRC using one of the pre-OMC pickoff RFPDs. At the sites, this auxiliary beam is provided by the newly-installed squeezer laser. Prior to injection, an AM sideband is imprinted on the auxiliary beam using an AOM and polarizer. The sinusoidal AOM drive signal is provided by a network analyzer, which sweeps in frequency across the MHz band and demodulates the PD signal in-phase to make an RF transfer function measurement. At the FSR, there will be a AM transmission resonance (reflection minimum). If HOMs are also present (created by either partially occluding or misaligning the injection beam), they too will generate transmission resonances, but at a frequency shift proportional to the Gouy phase. For the theoretical 19 deg one-way Gouy phase at the sites, this mode spacing is approximately 300 kHz. If the transmission resonances of two or more modes can be simultaneously measured, their frequency separation will provide a direct measurement of the SRC Gouy phase.

The above figure illustrates this measurement configuration. An attached PDF gives more detail and the expected response based on Finesse modeling of this IFO configuration.

  13716   Wed Mar 28 21:47:37 2018 arijitUpdateIOOMCREFL_PD Optical response measurement

Kevin, Gautam and Arijit

We did a optical measurement of the MCREFL_PD transimpedance using the Jenny Laser set-up. We used 0.56mW @1064nm on the NewFocus 1611 Photodiode as reference and 0.475mW @1064nm on the MCREFL_PD. Transfer function was measured using the AG4395 network analyzer. We also fit the data using the refined LISO model. From the optical measurement, we can see that we do not have a prominent peak at about 300MHz like the one we had from the electrical transimpedence measurement. We also put in the electrical transimpedence measurement as reference. RMS contribution of 300MHz peak to follow.

 

As per Rana`s advice I have updated the entry with information on the LISO fit quality and parameters used. I have put all the relevant files concerning the above measurement as well as the LISO fit and output files as a zip file "LISO_fit" . I also added a note describing what each file represents. I have also updated the plot with fit parameters and errors as in elog 10406.

ELOG V3.1.3-